From a basket of 10 pears and 7 peaches, a shopper must select 4 pears and 3 peaches. In how many ways can this be done?
Answer by jim_thompson5910(35256)

There are 10 C 4 = (10!)/(4!*(10-4)!) = 210 ways to select 4 pears (from the pool of 10). I'm using a combination since order doesn't matter.


There are 7 C 3 = (7!)/(3!*(7-3)!) = 35 ways to select 3 peaches (from the pool of 7). Again, we're using a combination because order doesn't matter.


So there are 210*35 = 7,350 ways to select 4 pears and 3 peaches (from a basket of 10 pears and 7 peaches).


Note: let me know if you need me to go over the combination formula.